# Bosenovae with Quadratically-Coupled Scalars in Quantum Sensing Experiments

### v1

@@ VI Conclusions
We observe that phenomenon of screening can significantly diminish ULDM detection prospects for sizable quadratic couplings to SM. This motivates further developments of space-based experiments that will avoid screening associated with the Earth affecting terrestrial experiments. Further, our analysis highlights that there is a ULDM parameter gap where there are no existing or planned space-based experiments that are sensitive to the larger mass range of $m_{\phi}\gtrsim 10^{-10}~{}{\rm eV}$.
@@ Acknowledgements
We thank Abhishek Banerjee, Gilad Perez, and Yevgeny Stadnik for useful discussions. This work was supported in part by the NSF QLCI Award OMA - 2016244, NSF Grant PHY-2012068, and the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (Grant Number 856415). JA thanks the International Center for Quantum-field Measurement Systems for Studies of the Universe and Particles (QUP) for the warm hospitality and exciting research environment. The work of JE and VT was supported by the World Premier International Research Center Initiative (WPI), MEXT, Japan, and by the JSPS KAKENHI Grant Numbers 21H05451 (JE), 21K20366 (JE) and 23K13109 (VT). The work of JE was also supported by the Swedish Research Council (VR) under grants 2018-03641 and 2019-02337. This research was supported in part by the INT’s U.S. Department of Energy grant No. DE-FG02- 00ER41132. This article is also based upon work from COST Action COSMIC WISPers CA21106, supported by COST (European Cooperation in Science and Technology).
@@ Appendix A Results with Superradiance Bounds
In the main text, our benchmark choice $\lambda=\lambda_{b}$ in Eq. (12) did not account for the existence of black hole superradiance (BHSR) constraints from rapidly-rotating black holes, discussed in Sec. II.3. In the presence of BHSR limits, we can instead use a discontinuous benchmark of the form
@@
Figure 8: Current and projected bounds on $d_{m_{e}}^{(2)}$ from terrestrial and space-based experiments as a function of $E_{\phi}$ including BHSR in $\lambda$ benchmark given in Eq. (30). The strongest limits for each mass are depicted by the solid black lines. The dashed and dotted black lines indicate the potential reach for detection of bosenovae at distances of $r=$ pc and kpc from the Earth, respectively. See Fig. 5 and Sec. IV for details and references of experimental lines.
Figure 9: Current and projected bounds on $d_{e}^{(2)}$ from terrestrial and space-based experiments as a function of $E_{\phi}$ including BHSR in $\lambda$ benchmark given in Eq. (30). The strongest limits for each mass are depicted by the solid black lines. The dashed and dotted black lines indicate the potential reach for detection of bosenovae at distances of $r=$ pc and kpc from the Earth, respectively. See Fig. 6 and Sec. IV for details and references of experimental lines.
Figure 10: Current and projected bounds on $d_{g}^{(2)}$ from terrestrial and space-based experiments as a function of $E_{\phi}$ including BHSR in $\lambda$ benchmark given in Eq. (30). The strongest limits for each mass are depicted by the solid black lines. The dashed and dotted black lines indicate the potential reach for detection of bosenovae at distances of $r=$ pc and kpc from the Earth, respectively. See Fig. 7 and Sec. IV for details and references of experimental lines.

### v2

@@ 6 Conclusions
We observe that the phenomenon of screening can significantly diminish ULDM detection prospects for sizable quadratic couplings to SM. This motivates further developments of space-based experiments that will avoid screening associated with the Earth affecting terrestrial experiments. Further, our analysis highlights that there is a ULDM parameter gap where there are no existing or planned space-based experiments that are sensitive to the larger mass range of $m_{\phi}\gtrsim 10^{-10}~{}{\rm eV}$.
@@ Acknowledgements
We thank Abhishek Banerjee, Gilad Perez, and Yevgeny Stadnik for useful discussions. This work was supported in part by the NSF QLCI Award OMA - 2016244, NSF Grants PHY-2012068, PHY-2309254, and the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (Grant Number 856415). JA thanks the International Center for Quantum-field Measurement Systems for Studies of the Universe and Particles (QUP) for the warm hospitality and exciting research environment. The work of JE and VT was supported by the World Premier International Research Center Initiative (WPI), MEXT, Japan, and by the JSPS KAKENHI Grant Numbers 21H05451 (JE), 21K20366 (JE) and 23K13109 (VT). The work of JE was also supported by the Swedish Research Council (VR) under grants 2018-03641 and 2019-02337. This research was supported in part by the INT’s U.S. Department of Energy grant No. DE-FG02- 00ER41132. This article is also based upon work from COST Action COSMIC WISPers CA21106, supported by COST (European Cooperation in Science and Technology).
@@ Appendix A Results with Superradiance Bounds
In the main text, our benchmark choice $\lambda=\lambda_{b}$ in Eq. (12) did not account for the existence of black hole superradiance (BHSR) constraints from rapidly-rotating black holes, discussed in Sec. 2.3. In the presence of BHSR limits, we can instead use a discontinuous benchmark of the form
@@
Figure 8: Current and projected bounds on $d_{m_{e}}^{(2)}$ from terrestrial and space-based experiments as a function of $E_{\phi}$ including BHSR in $\lambda$ benchmark given in Eq. (30). The strongest limits for each mass are depicted by the solid black lines. The dashed and dotted black lines indicate the potential reach for detection of bosenovae at distances of $r=$ pc and kpc from the Earth, respectively. See Fig. 5 and Sec. 4 for details and references of experimental lines.
Figure 9: Current and projected bounds on $d_{e}^{(2)}$ from terrestrial and space-based experiments as a function of $E_{\phi}$ including BHSR in $\lambda$ benchmark given in Eq. (30). The strongest limits for each mass are depicted by the solid black lines. The dashed and dotted black lines indicate the potential reach for detection of bosenovae at distances of $r=$ pc and kpc from the Earth, respectively. See Fig. 6 and Sec. 4 for details and references of experimental lines.
Figure 10: Current and projected bounds on $d_{g}^{(2)}$ from terrestrial and space-based experiments as a function of $E_{\phi}$ including BHSR in $\lambda$ benchmark given in Eq. (30). The strongest limits for each mass are depicted by the solid black lines. The dashed and dotted black lines indicate the potential reach for detection of bosenovae at distances of $r=$ pc and kpc from the Earth, respectively. See Fig. 7 and Sec. 4 for details and references of experimental lines.